What does the x's equal? -10=1/2x+x
Answer by checkley77(12844)

-10=1/2x+x
-10=X/2+X
-10=(X+2X)/2
-10*2=3X
3X=-20
X=-20/3 ANS.
PROOF:
-10=(-20/3)/2-20/3
-10=-20/3*1/2-20/3
-10=-20/6-20/3
-10=(-20-20*2)/6
-10=(-20-40)/6
-10=-60/6
-10=-10